# Die Hard: The On-Off-Cycle of Galaxies on the Star Formation Main Sequence

### v1

@@ 3.1 Formation History Classification
We want to understand if galaxies live on the main sequence until quenching or not. Thus, we trace galaxies across cosmic time and classify them according to their star formation rate evolution relative to the star formation main sequence by Pearson et al. (2018). While it is evident from Fig. 1 that the relation found by Pearson et al. (2018) does not fit the best to the Magneticum galaxies, we chose to use their best-fit relation as a reference as they provide a consistent method in obtaining the main sequence covering a large range of redshifts, which is required to study the long-term evolution of galaxies. In addition, their values are based on the same cosmology from WMAP7 results (Komatsu et al. 2011) as well as the IMF by Chabrier (2003). However, we do not limit the study to the relation found by Pearson et al. (2018) but rather explore other main sequence definitions as baseline as well, which we will mention throughout this work if differences to the results based on the main sequence by Pearson et al. (2018) are found. Overall, the results presented in this work do not change with the choice of base main sequences, and only the relative fractions of properties vary depending on that choice. An overview of the main sequences used as references and the obtained results is given in Tab. 1. For each snapshot, we define each galaxy as:
  • $\bullet$
    quiescent if it is more than 0.5 dex below the main sequence value $\mathrm{SFR_{\mathrm{MS}}(M_{*},z)}$, reaching down to SFR$=0$,
  • $\bullet$
    star forming if it is within a ±$\pm 0.5$ dex range of the main sequence value,
  • $\bullet$
@@
Figure 2: Examples of the four formation scenarios for galaxies with respect to specific star formation rate changes over time. The black lines and shades indicate the main sequence values depending on the redshift and stellar mass of a galaxy according to results by Pearson et al. (2018) with a scatter of ±$\pm 0.5\ \mathrm{dex}$ applied to differentiate between on or off the main sequence. The panels show a main-sequence-type galaxy (cyan), an early-quenched galaxy (dark red), a rejuvenating galaxy (magenta) and a caught galaxy (blue) from top to bottom. Two classes are not explicitly represented here. First, the “late-quenched” types only differ from the “early-quenched” by dropping off the main sequence after redshift $z=1$. Second, rejuvenator-main-sequence-type galaxies recover a stable position on the main sequence at redshift $z=0$, unlike rejuvenator-quenched-type galaxies which also go through rejuvenation cycles but do not end up on the main sequence. The vertical lines highlight the snapshots corresponding to each row in Figure 3.
@@
The study was performed using the star-forming main sequence fit by Pearson et al. (2018). However, we have reproduced our classifications using also other fits from observations as shown in Table 1, as well as the quiescence limits by Franx et al. (2008); Carnall et al. (2020). While the absolute numbers of galaxies in each class change, a consistent picture emerges, with only a very small fraction of galaxies evolving along the star-forming main sequences. The bulk of galaxies quenches at high redshift and a consistently significant fraction of galaxies hops on and off the main sequence, of which also a large part end up on the main sequence at $z\approx 0$. As a result, we find that the dominant group of galaxies on the main sequence at $z\approx 0$ have experienced passive phases on gigayear time scales, independent of the exact main sequence definition used for classification. Consequently, our classifications of these different star formation behaviors are robust and not simply definition dependent. We do note, however, a systematic difference between the main sequence results depending on wether or not a turnover mass was included in their fit. Those main sequence fits which do (Leslie et al. 2020; Popesso et al. 2023) result in an overall greater number of rejuvenated galaxies (‘RM.’ in Table 1) and less galaxies that always remained on the main sequence (MS) as compared to those which do not include a turnover mass (Speagle et al. 2014; Pearson et al. 2018). This is expected due to the redshift evolution of each main sequence description. While the relations by Leslie et al. (2020); Popesso et al. (2023) require higher star formation rates at high redshift than the relations by Speagle et al. (2014); Pearson et al. (2018), it is the other way round towards $z=0$. In addition, the turnover mass amplifies this effect for more massive galaxies, leading us to the conclusion that rejuvenation is extremely common among galaxies on main sequences with turnover mass.
[…]
@@
Table 1: Numbers of galaxies within each formation class from top to bottom: “main sequence”, “quenched”, “rejuvenating” and “caught” with respect to expectations according to a list of results for the star forming main sequence by Speagle et al. (2014); Pearson et al. (2018); Leslie et al. (2020); Popesso et al. (2023). The active-quiescent limits applied in the work by Franx et al. (2008); Carnall et al. (2020) were also used for comparison with an approach different from the SFMS. Caught-type galaxies were manually identified, which was only done for the first two references listed. Sub-classes of quenched and rejuvenating galaxies are listed in the four lower rows with “quenched early” (i.e. at $z>1$), “quenched late” (i.e. at $z<1$), “rejuvenating quenched” and “rejuvenating main sequence” galaxies. Note that the total number of galaxies varies due to the sample selection criteria as discussed in section 2.
In Figure 3 we plot the star formation rate versus stellar mass for galaxies of each class (columns, going from left to right as main sequence, quenched, rejuvenator and caught) from $z=5$ down to $z=0.1$ (rows, top to bottom). A galaxy that has not yet fallen off the main sequence is plotted in light blue, and galaxies that have fallen off of the sequence once, twice or three times are shown in red, orange and yellow. The thin black contour lines show the distribution of our full galaxy sample at each redshift for comparison. We also show the main sequence as defined by Pearson et al. (2018) as a black line (with the shaded area representing ±$\pm 0.5$ dex).
As noted, we find the main bulk of our galaxies to be either quenched or rejuvenators. The former show a large spread in quenching times, with some falling off the sequence as early as $z=2.3$, while others remain on the main sequence all the way down to $z=0.5$ before finally quenching. This is true also for the rejuvenators, except that they continue returning back up to the main sequence. We find that most rejuvenator galaxies undergo their first quenching process at $12>z>$ 1 just after the cosmic peak of star formation (Behroozi et al. 2013; Madau & Dickinson 2014).
In the following, we will discuss each of the classes separately in more detail with the main sequence by Pearson et al. (2018) as a reference.
Figure 3: Star formation main sequence evolution of the Magneticum galaxies traced back from the selection made at z=0 (bottom-row panels). The rows contain snapshots with decreasing redshift from top to bottom between $06<z<$ 0. From left to right, the different sub-samples “Main Sequence”, “Quenched”, “Rejuvenator” and “Caught” are highlighted by colors in front of the total sample in grey in the background. Light blue data points represent galaxies that have not yet dropped below the main sequence; each quenching event is counted and encoded in the changing color of the data points as red (once), orange (twice) and yellow (three times). The large points are the cherry-picked representative galaxies from Figure 2. Quenched representative galaxies are set as half-circles to the bottom of the panels for visibility. The black line and grey shades correspond the main-sequence found by Pearson et al. (2018). Grey contours outline the distribution of the full galaxy sample for comparison in each panel.
@@ 3.3 Quenched Galaxies
Figure 4: Similar to Figure 2 but focused on the difference between the subclass pairs of early- and late-quenched galaxies as well as quenched and main-sequence rejuvenators. The colored dashed lines represent the cases also shown in Figure 2, while the colored solid lines are the subtypes to differentiate. The grey dashed quiescence criteria (Carnall et al. 2020)) and black-shaded SFMS expectations (Pearson et al. 2018) are derived from the stellar masses of the latter. This causes the pink dashed main-sequence-rejuvenator to not lie on the SFMS at $z\approx 0$ due to the larger stellar mass (see Figure 5) and therefore lower sSFR. The early-quenched galaxy does recover star-formation after quenching but never consistently makes it back onto the main sequence. The late-quenched galaxy also almost makes a passive rejuvenation akin to caught-type galaxies, since the main sequence drops at $z<0.5$. However, it is still mostly more than 0.5 dex below the main sequence averaged over a smoothing window of one gigayear, ruling out a rejuvenation scenario. The main-sequence rejuvenator undergoes a more pronounced quenched phase than the quenched rejuvenator.
@@
More and less rapid quenching scenarios are found among galaxies of both sub-classes, as well as small and short-lived rekindling of star formation. To illustrate this distinction further, we plot in the left panel of Figure 5 the star formation rate against stellar mass of galaxies from the ”quenched” class at five different redshifts as in Figure 3, but change the color scheme. Galaxies classified as early quenched are always plotted in red, while late quenchers are always in blue. This allows us to see that the majority of early quenchers still lie on or near the main sequence at $z=2.3$, such that their quenching must occur abruptly between $12>z>$ 1. Conversely, late quenchers do not surpass the mass range of the figure at $z=2.3$ and moment of quenching is less temporally confined, with occurrences spreading evenly between the time steps $0.51>z>$ 0.5 and $00.5>z>$ 0. This hints at different causes of quenching between early and late quenching, such as feedback and outflows as opposed to starvation or merging.
@@
Figure 5: Similar to Figure 3 but focused on the difference between the subclass pairs of early- and late-quenched galaxies as well as quenched and main-sequence rejuvenators. In this version, the colors refer to the subclass types instead of quenching incidents. The large points show the cherry picked representative galaxies from Figure 4, again set to the bottom of the panels as half-circles for visibility in case of low SFR values. By construction, the difference between early- and late-quenched is most pronounced at $z=1$. At $z\approx 0$, the non-zero distributions are barely noticeable. Analogously, the difference between the two rejuvenator subclasses is the sharpest at $z\approx 0$. Both subclasses undergo rejuvnenation cycles at different times, but the relation to the SFMS determines the classification.
@@ 3.4 Rejuvenating Galaxies
Figure 6: Median radial profiles and 68th-percentile spread of stellar ages at $z=0$. Before extracting the running median and scatter, we calculated the mean stellar ages for each galaxy in nine equal-width radial bins between 00 and $5R_{\mathrm{1/2}}$. The solid lines and shaded regions correspond to the statistics of the class indicated at the top of each panel. The dashed lines are the the running medians from the panels to the left of each panel for a direct comparison. Higher values correspond to an older population, which is fairly constant at around ten gigayears for early-quenched galaxies. Main-sequence-type galaxies show increasingly young populations towards the center, while the three types of rejuvenating classes are the youngest at roughly one half-mass radius and increase in age again towards the very center as well as the outskirts.
To further study the composition of active galaxies at $z\approx 0$, we differentiate rejuvenating galaxies that end up on the main sequence (in this work referred to as “Rejuvenator Main Sequence” or “RM”) from those that end up quenched (“Rejuvenator Quenched” or “RQ”). The conceptual difference is visualized with the chosen examples in the lower panel of Figure 4, where we plot the sSFR against cosmic time of two example rejuvenator galaxies, one which ends up quenched (RQ, solid lilac line) and one which ends up star forming (RM, dashed pink line). Even though the quenched rejuvenator has a more consistent SFH, it drops off into a new quenched phase towards $z\approx 0$. The main-sequence rejuvenator undergoes an extended period of quiescence until rejuvenating back onto the main sequence in two rejuvenation cycles, ending up as actively star forming. In the right-hand column of Figure 5, we can see that the two subclasses of rejuvenators are not distinguishable by their location in the star formation rate versus stellar mass plot at higher redshifts, instead lying alternatively on or off the main sequence. We therefore conclude that there is no physical difference between the two classes of rejuvenators (unlike for the early and late quenchers), with the only distinguishing feature being the timing of their rejuvenation.
@@ 3.5 Caught Galaxies
A rare but nonetheless peculiar class of galaxies, namely $\sim 2\%$ of our sample, consists of such whose star formation rates drop below the main sequence at some point but never fully cease their star formation. Instead, they very continuously form stars with a nearly constant rate just below the main sequence, until they reappear on it once the overall main sequence itself has lowered to reach the SFR of these galaxies. An example case of this formation scenario is shown in the bottom panel of Figure 2. We distinguish these galaxies from the rejuvenators, as they never truly rise in their SFR (i.e., rejuvenate), and instead are a result of the overall decline in cosmic SFRD. Members of this galaxy class maintain a constant level of lower star formation rate, such that caught-type galaxies are especially promising candidates to test the equilibrium prediction of the “Bathtub” model (Burkert 2017). Therein, a constant gas accretion rate leads to an equilibrium state between accretion, star formation and feedback, resulting in a constant gas mass and star formation rate. This formation scenario is particularly interesting since they are technically more stably retaining their star formation compared to the average main sequence galaxy, which is actually lowering its SFR with redshifts. Their small sample size is expected as they continuously contribute to the scatter of the main sequence opposing the definition of a well-constrained star formation main sequence.

### v2

@@ 3.1 Formation history classification
We want to understand if galaxies live on the main sequence until quenching or not. Thus, we trace galaxies across cosmic time and classify them according to their star formation rate evolution relative to the star formation main sequence by pearson:2018. While it is evident from Fig. 1 that the relation found by pearson:2018 does not fit the best to the Magneticum galaxies, we chose to use their best-fit relation as a reference as they provide a consistent method in obtaining the main sequence covering a large range of redshifts, which is required to study the long-term evolution of galaxies. In addition, their values are based on the same cosmology from WMAP7 results (komatsu:2011) as well as the IMF by chabrier:2003. However, we do not limit the study to the relation found by pearson:2018 but rather explore other main sequence definitions as baseline as well, which we will mention throughout this work if differences to the results based on the main sequence by pearson:2018 are found. Overall, the results presented in this work do not change with the choice of base main sequences, and only the relative fractions of properties vary depending on that choice. An overview of the main sequences used as references and the obtained results is given in Tab. 2. For each snapshot, we define each galaxy as:
    quiescent if it is more than 0.5 dex below the main sequence value $\mathrm{SFR_{\mathrm{MS}}(M_{*},z)}$, reaching down to SFR=0=0,
    star forming if it is within a $\pm 0.5$ dex range of the main sequence value,
@@
Figure 2: Examples of the four formation scenarios for galaxies with respect to specific star formation rate changes over time. The black lines and shades indicate the main sequence values depending on the redshift and stellar mass of a galaxy according to results by pearson:2018 with a scatter of $\pm 0.5\ \mathrm{dex}$ applied to differentiate between on or off the main sequence. The panels show a main-sequence-type galaxy (cyan), an early-quenched galaxy (dark red), a rejuvenating galaxy (magenta) and a caught galaxy (blue) from top to bottom. Two classes are not explicitly represented here. First, the “late-quenched” types only differ from the “early-quenched” by dropping off the main sequence after redshift $z=1$. Second, rejuvenator-main-sequence-type galaxies recover a stable position on the main sequence at redshift $z=0$, unlike rejuvenator-quenched-type galaxies which also go through rejuvenation cycles but do not end up on the main sequence. The vertical lines highlight the snapshots corresponding to each row in Figure 3.
@@
The study was performed using the star-forming main sequence fit by pearson:2018. However, we have reproduced our classifications using also other fits from observations as shown in footnote 2, as well as the quiescence limits by franx:2008; carnall:2020. While the absolute numbers of galaxies in each class change, a consistent picture emerges, with only a very small fraction of galaxies evolving along the star-forming main sequences. The bulk of galaxies quenches at high redshift and a consistently significant fraction of galaxies hops on and off the main sequence, of which also a large part end up on the main sequence at $z\approx 0$. As a result, we find that the dominant group of galaxies on the main sequence at $z\approx 0$ have experienced passive phases on gigayear time scales, independent of the exact main sequence definition used for classification. Consequently, our classifications of these different star formation behaviors are robust and not simply definition dependent. We do note, however, a systematic difference between the main sequence results depending on wether or not a turnover mass was included in their fit. Those main sequence fits which do (leslie:2020; popesso:2023) result in an overall greater number of rejuvenated galaxies (‘RM.’ in footnote 2) and less galaxies that always remained on the main sequence (MS) as compared to those which do not include a turnover mass (speagle:2014; pearson:2018). This is expected due to the redshift evolution of each main sequence description. While the relations by leslie:2020; popesso:2023 require higher star formation rates at high redshift than the relations by speagle:2014; pearson:2018, it is the other way round towards $z=0$. In addition, the turnover mass amplifies this effect for more massive galaxies, leading us to the conclusion that rejuvenation is extremely common among galaxies on main sequences with turnover mass.
Table 1: Numbers of galaxies within each formation class with respect to expectations according to a list of results for the SFMS.  [^2] Classes from top to bottom: “main sequence”, “quenched”, “rejuvenating” and “caught”. References from left to right: pearson:2018; speagle:2014; leslie:2020; popesso:2023. The active-quiescent limits applied in the work by franx:2008; carnall:2020 were also used for comparison with an approach different from the SFMS. Caught-type galaxies were manually identified, which was only done for the first two references listed. Subclasses of quenched and rejuvenating galaxies are listed in the four lower rows with “quenched early” (i.e. at $z>1$), “quenched late” (i.e. at $z<1$), “rejuvenating quenched” and “rejuvenating main sequence” galaxies. Note that the total number of galaxies varies due to the sample selection criteria as discussed in section 2.
@@
In Figure 3 we plot the star formation rate versus stellar mass for galaxies of each class (columns, going from left to right as main sequence, quenched, rejuvenator and caught) from $z=5$ down to $z=0.1$ (rows, top to bottom). The circle colors in this figure indicate the number of transitions a galaxy has undergone from star-forming to passive. A galaxy that has not yet fallen off the main sequence by the respective redshift is shown in light blue. Galaxies that have quenched once appear as red circles, so a red circle on the main sequence has rejuvenated. Orange colors indicate a second transition into quiescence after rejuvenation and yellow colors indicate a third drop off the SFMS after two rejuvenation cycles. The thin black contour lines show the distribution of our full galaxy sample at each redshift for comparison. We also show the main sequence as defined by pearson:2018 as a black line (with the shaded area representing $\pm 0.5$ dex).
As noted, we find the main bulk of our galaxies to be either quenched or rejuvenators. The former show a large spread in quenching times, with some falling off the sequence as early as $z=2.3$, while others remain on the main sequence all the way down to $z=0.5$ before finally quenching. This is true also for the rejuvenators, except that they continue returning back up to the main sequence. We find that most rejuvenator galaxies undergo their first quenching process at $2>z>1$ just after the cosmic peak of star formation (behroozi:2013; madau:2014:dickinson).
In the following, we will discuss each of the classes separately in more detail with the main sequence by pearson:2018 as a reference.
Figure 3: Star formation main sequence evolution of the Magneticum galaxies traced back from the selection made at z=0 (bottom-row panels). The rows contain snapshots with decreasing redshift from top to bottom between $6<z<0$. From left to right, the different sub-samples “Main Sequence”, “Quenched”, “Rejuvenator” and “Caught” are highlighted by colors in front of the total sample in gray in the background. Light blue data points represent galaxies that have not yet dropped below the main sequence; each quenching event is counted and encoded in the changing color of the data points as red (once), orange (twice) and yellow (three times). The large points are the cherry-picked representative galaxies from Figure 2. Quenched representative galaxies are set as half-circles to the bottom of the panels for visibility. The black line and gray shades correspond the main-sequence found by pearson:2018. Gray contours outline the distribution of the full galaxy sample for comparison in each panel.
@@ 3.3 Quenched galaxies
Figure 4: Similar to Figure 2 but focused on the difference between the subclass pairs of early- and late-quenched galaxies as well as quenched and main-sequence rejuvenators. The colored dashed lines represent the cases also shown in Figure 2, while the colored solid lines are the subtypes to differentiate. The gray dashed quiescence criteria (carnall:2020)) and black-shaded SFMS expectations (pearson:2018) are derived from the stellar masses of the latter. This causes the pink dashed main-sequence-rejuvenator to not lie on the SFMS at $z\approx 0$ due to the larger stellar mass (see Figure 5) and therefore lower sSFR. The early-quenched galaxy does recover star-formation after quenching but never consistently makes it back onto the main sequence. The late-quenched galaxy also almost makes a passive rejuvenation akin to caught-type galaxies, since the main sequence drops at $z<0.5$. However, it is still mostly more than 0.5 dex below the main sequence averaged over a smoothing window of one gigayear, ruling out a rejuvenation scenario. The main-sequence rejuvenator undergoes a more pronounced quenched phase than the quenched rejuvenator.
@@
More and less rapid quenching scenarios are found among galaxies of both subclasses, as well as small and short-lived rekindling of star formation. To illustrate this distinction further, we plot in the left panel of Figure 5 the star formation rate against stellar mass of galaxies from the ”quenched” class at five different redshifts as in Figure 3, but change the color scheme. Galaxies classified as early quenched are always plotted in red, while late quenchers are always in blue. This allows us to see that the majority of early quenchers still lie on or near the main sequence at $z=2.3$, such that their quenching must occur abruptly between $2>z>1$. Conversely, late quenchers do not surpass the mass range of the figure at $z=2.3$ and moment of quenching is less temporally confined, with occurrences spreading evenly between the time steps $1>z>0.5$ and $0.5>z>0$. This hints at different causes of quenching between early and late quenching, such as feedback and outflows as opposed to starvation or merger accretion.
The median values of net gas accretion rates relative to dark matter assembly in Figure 13 in the appendix support the image of an early starburst with a rapid transition toward net outflows for early-quenched galaxies. The relation between black hole mass and stellar mass shown in Figure 14 in the appendix implies a correlation with AGN feedback, as early-quenched galaxies have the most massive black holes. As shown by kimmig:2025a, an early and rapid formation also comes with an efficient gas transport towards the AGN. In addition to the starburst the feedback by the massive black hole then adds to the complete removal of gas leading to a permanent quenching. Late-quenched galaxies have lower accretion rate peaks and black hole masses, indicating a less rapid evolution due to gas accretion. The prevalence and impact of mergers in the two quenching scenarios will be disentangled in future work.
@@
Figure 5: Similar to Figure 3 but focused on the difference between the subclass pairs of early- and late-quenched galaxies as well as quenched and main-sequence rejuvenators. In this version, the colors refer to the subclass types instead of quenching incidents. The large points show the cherry picked representative galaxies from Figure 4, again set to the bottom of the panels as half-circles for visibility in case of low SFR values. By construction, the difference between early- and late-quenched is most pronounced at $z=1$. At $z\approx 0$, the non-zero distributions are barely noticeable. Analogously, the difference between the two rejuvenator subclasses is the sharpest at $z\approx 0$. Both subclasses undergo rejuvenation cycles at different times, but the relation to the SFMS determines the classification.
@@ 3.4 Rejuvenating galaxies
Figure 6: Median radial profiles and 68th-percentile spread of stellar ages at $z=0$. Before extracting the running median and scatter, we calculated the mean stellar ages for each galaxy in nine equal-width radial bins between 0 and $5R_{\mathrm{1/2}}$. The solid lines and shaded regions correspond to the statistics of the class indicated at the top of each panel. The dashed lines are the the running medians from the panels to the left of each panel for a direct comparison. Higher values correspond to an older population, which is fairly constant at around ten gigayears for early-quenched galaxies. Main-sequence-type galaxies show increasingly young populations towards the center, while the three types of rejuvenating classes are the youngest at roughly one half-mass radius and increase in age again towards the very center as well as the outskirts.
To further study the composition of active galaxies at $z\approx 0$, we differentiate rejuvenating galaxies that end up on the main sequence (in this work referred to as “Rejuvenator Main Sequence” or “RM”) from those that end up quenched (“Rejuvenator Quenched” or “RQ”). The conceptual difference is visualized with the chosen examples in the lower panel of Figure 4, where we plot the sSFR against cosmic time of two example rejuvenator galaxies, one which ends up quenched (RQ, solid lilac line) and one which ends up star forming (RM, dashed pink line). Even though the quenched rejuvenator has a more consistent SFH, it drops off into a new quenched phase towards $z\approx 0$. The main-sequence rejuvenator undergoes an extended period of quiescence until rejuvenating back onto the main sequence in two rejuvenation cycles, ending up as actively star forming. In the right-hand column of Figure 5, we can see that the two subclasses of rejuvenators are not distinguishable by their location in the star formation rate versus stellar mass plot at higher redshifts, instead lying alternatively on or off the main sequence. We therefore conclude that there is no physical difference between the two classes of rejuvenators (unlike for the early and late quenchers), with the only distinguishing feature being the timing of their rejuvenation. Intriguingly, we find that rejuvenating galaxies host less massive black holes (see Figure 14 in the appendix), which may facilitate rejuvenation with less feedback emitted. We note also that the impact of mergers will be studied in future work using merger trees, which connect SUBFIND-identified progenitors of a galaxy through the simulation snapshots. This will allow us to look for differences between galaxy evolution classes regarding the time correlation between merger events and quenching or rejuvenation.
@@ 3.5 Caught galaxies
A rare but nonetheless peculiar class of galaxies, namely $\sim 2\%$ of our sample, consists of such whose star formation rates drop below the main sequence at some point but never fully cease their star formation. Instead, they very continuously form stars with a nearly constant rate just below the main sequence, until they reappear on it once the overall main sequence itself has lowered to reach the SFR of these galaxies. An example case of this formation scenario is shown in the bottom panel of Figure 2. We distinguish these galaxies from the rejuvenators, as they never truly rise in their SFR (i.e., rejuvenate), and instead are a result of the overall decline in cosmic SFRD. Members of this galaxy class maintain a constant level of lower star formation rate, such that caught-type galaxies are especially promising candidates to test the equilibrium prediction of the “Bathtub” model (burkert:2017). Therein, a constant gas accretion rate leads to an equilibrium state between accretion, star formation and feedback, resulting in a constant gas mass and star formation rate. This formation scenario is particularly interesting since they are technically more stably retaining their star formation compared to the average main sequence galaxy, which is actually lowering its SFR with redshifts. Their small sample size is expected as they continuously contribute to the scatter of the main sequence opposing the definition of a well-constrained star formation main sequence.